# AffordanceSAM: Segment Anything Once More in Affordance Grounding

### v1

@@ Abstract
Improving the generalization ability of an affordance grounding model to recognize regions for unseen objects and affordance functions is crucial for real-world application. However, current models are still far away from such standards. To address this problem, we introduce AffordanceSAM, an effective approach that extends SAM’s generalization capacity to the domain of affordance grounding. For the purpose of thoroughly transferring SAM’s robust performance in segmentation to affordance, we initially propose an affordance-adaption module in order to help modify SAM’s segmentation output to be adapted to the specific functional regions required for affordance grounding. We concurrently make a coarse-to-fine training recipe to make SAM first be aware of affordance objects and actions coarsely, and then be able to generate affordance heatmaps finely. Both quantitative and qualitative experiments show the strong generalization capacity of our AffordanceSAM, which not only surpasses previous methods under AGD20K benchmark but also shows evidence to handle the task with novel objects and affordance functions.
1 Introduction
@@
Affordance grounding [12, 15] refers to finding potential “action possibilities” regions of an object, which plays a pivotal role in bridging the gap between visual perception and robotic action. Recently, attempts [29, 24, 39, 25, 34, 38] made to endow models to have the ability to ground objects can be broadly divided into two methodologies. The first type of methods is weakly supervised affordance grounding [29, 24, 34], which aims to make the right classifications on the predefined affordance classes by extracting affordance-related knowledge from different centric views. The affordance maps are obtained via class activation mapping (CAM) [56]. However, these methods completely break down when facing new affordance functions as their supported classes are fixed. To tackle this problem, the second type of methods uses affordance maps to directly supervise the models [39, 25, 38]. However, the brittle affordance components trained from scratch on the hundreds of pieces manually labeled training data make it insufficient to obtain a highly generalized model. To this end, we raise a question: How can we obtain a model with generalization ability under such limited data conditions?
Figure 1: Visualization of differences between segmentation and affordance, where segmentation focuses on separating objects directly according to the prompt, but affordance emphasizes grounding the possible effective part of objects based on the affordance query.
In this work, we try to decompose and answer the above question in a step-by-step manner:
@@
(i) Can we incorporate a suitable and generalized vision prior? With the recent advancements in large-scale pre-trained foundation models, many works have attempted to leverage the prior knowledge for downstream task transfer [53, 47, 14, 18]. We believe this paradigm is also effective for affordance grounding task. Given that affordance grounding inherently requires pixel-level localization, we explore the possibility of incorporating SAM (Segment Anything Model) [20, 41] since SAM not only demonstrates remarkable generalization capabilities but also excels in precise object localization after pre-trained on large-scale masked-labeled data. Moreover, the model structure and decoder output format of segmentation and affordance grounding are highly similar. Considering these similarities between segmentation and affordance grounding, we hypothesize that incorporating SAM could be both promising and well-suited for our problem.
@@
(ii) How can we effectively utilize and transfer SAM’s comprehensive knowledge into affordance grounding task? Although there are some similarities between segmentation and affordance grounding, task differences still exist. As shown in Figure 1, firstly, affordance grounding requires the recognition of functional parts of objects based on verb queries, while segmentation only requires the separation of complete objects; next, the ground-truth of segmentation and affordance differs slightly: the former uses a binary mask, while the latter uses a heatmap that encodes functional possibilities. To address these two primary challenges, we propose a simple but efficient framework from model structure to specific training recipe to obtain an affordance version of SAM, termed as AffordanceSAM.
@@
Specifically, we initially design an affordance-adaption module that utilizes learnable queries to interact with text and image features, these queries are then sent into the decoder to refine the original mask produced by SAM. We assume that this prompt learning-like [57] lightweight module can effectively search for affordance-relevant parts in the comprehensive knowledge of SAM. We then make a three-stage training recipe that employs a coarse-to-fine approach. In the first stage, we curate a coarsely mask-labeled dataset by merging and reclassifying existing datasets, focusing on object-affordance pairings. This enables the model to learn basic object-verb relationships. After that, we use a weakly supervised model to generate pseudo-labels for unannotated images and post-process them to supervise our model. Finally, we further fine-tune the model on the high-quality, human-labeled data for finer affordance map generation.
@@
  • We innovatively incorporate SAM for affordance grounding, empowering our model to generalize effectively beyond the training data.
  • We simultaneously introduce an affordance-adaption module and coarse-to-fine training recipe to effectively utilize and transfer SAM’s comprehensive knowledge into affordance grounding.
  • In our comprehensive experiments, our model achieves state-of-the-art (SOTA) performance under the AGD20K benchmark, and shows evidence to generalize well to novel actions and objects.
@@
Figure 2: (a). Architecture: AffordanceSAM is built upon EVF-SAM  [52] with an additional affordance-adaption module. The inputs of our model include a single image and a text prompt related to interaction. In our affordance-adaption module, we employ learnable queries to fuse with text feature and multiple layers of features from SAM encoder and then incorporate them into the original decoder to help the transformation from mask to the final affordance map. (b). Training: AffordanceSAM adopts a coarse-to-fine training recipe. In stage1 and stage3, we use ground truth binary masks and affordance maps to supervise the model, respectively. In stage2, affordance maps output by LOCATE and then pass through our proposed post-process program are used to supervise the model.
2 Related Work
We discuss with the most relevant studies here and provide a more discussion in Appendix 1.
Affordance grounding. Different from well-known detection [55] and segmentation [32] that acquire the model to output bounding box and binary mask of the desired objects, affordance grounding task [12, 15] needs model to output heatmap that encodes functional possibilities in a object. At the beginning, there are some works [34, 31, 36] tried to equip the model of such capacity. But because the data restriction at that time, these models can only recognize a few types of objects and actions. Next, Luo et al. collect the first large-scale affordance dataset called AGD20K [29] and annotatied a few of the data with affordance maps (1675/23816). After that, many works [24, 39, 25] engined by AGD20K have emerged. For example, LOCATE [24] localizes and extracts affordance-specific information in exocentric and transfers this knowledge to egocentric in a weakly supervised manner; OOAL [25] builds a vision-language learning framework based on CLIP [40] and DINOv2 [35] to conduct one-shot open affordance learning. However, non of above works shows satisfactory generalization ability to recognize regions for unseen objects and affordance functions (reasons are explained in Section 1 and evidences are shown in our experiment results). In this study, we aim at utilizing SAM as base model with our proposed transferring approach to build a generalized affordance grounding model to fill this gap.
Adapting foundation model to downstream task. Vision foundation models (e.g. SAM [20], CLIP [40], DINOv2 [35]) are large-sized models trained over a vast amounts of data . After training, they are suitable to be a start point for variety of downstream tasks under limited data condition. Recently, many works have attempted to leverage different foundation models for different downstream tasks [53, 47, 14, 18]. For example, ActionCLIP [47] proposes “pre-train, prompt and fine-tune” strategy to transfer CLIP to action recognition aspect; FM-FSOD [14] use DINOv2 to extract fine-grained features for few-shot object detection. Our work also share similarities, while we focus on adapting SAM to affordance grounding task with our proposed affordance-adaption module and coarse-to-fine training recipe.
3 Approach
The overall model architecture and training procedure of AffordanceSAM are shown in Figure 2(a) and Figure 2(b). As our model consists of two primary components: the EVF-SAM [52] [^1] EVF-SAM can be seen as an extension version of SAM that support text prompts. In this paper, we use term ‘EVF-SAM’ to refer the specific model part and ‘SAM’ for our general argument. baseline and the affordance-adaption module, we begin with describing each in Section 3.1 and Section 3.2. Finally, in Section 3.3, we provide a comprehensive elucidation of our coarse-to-fine training recipe.
3.1 Preliminary
Stage Data Source Label Type Num.
@@ 3 Labeled Part of AGD20K Human Annotated Affordance Map 1135 (Easy Split)
Table 1: Overview of the training datasets for AffordanceSAM. This table summarizes the data source, format of groundtruth, the number of samples for training in each stage. More details can be found in Appendix 2.
Baseline EVF-SAM. As we need to provide text prompts that contain affordance queries to the model but original SAM lacks language understanding abilities. We choose to start from EVF-SAM [52] which is the state-of-the-art method among attempts [52, 23, 42, 54] that explored SAM to segment objects according to the text instructions. We seek to leverage its robust segmentation capabilities for affordance grounding.
The EVF-SAM original framework incorporates four key components: a multimodal encoder (BEIT-3 [48]) denoted as $\mathcal{E}_{M}$, a prompt encoder $\mathcal{E}_{P}$, a SAM image encoder $\mathcal{E}_{I}$, and a SAM mask decoder $\mathcal{D}$. The input image is firstly processed through two parallel paths: (i): conversion to SAM image tokens $\mathbf{I}_{s}\in\mathbb{R}^{B\times N_{s}\times D_{s}}$, and (ii): transformation to multimodal image tokens $\mathbf{I}_{m}\in\mathbb{R}^{B\times N_{m}\times D_{m}}$, where $B$ represents batch size, $N$ sequence length, and $D$ feature dimension. Meanwhile, the text input is first tokenized to produce text tokens $\mathbf{T}\in\mathbb{R}^{B\times N_{t}\times D_{m}}$. These text and multimodal image tokens are then combined with a learnable [CLS] token $\in\mathbb{R}^{B\times 1\times D_{m}}$ and processed by the multimodal encoder:
@@
3.2 Affordance-Adaption Module
As shown in Figure 1, there exists a substantial disparity between the output of the segmentation task and the affordance grounding task. To efficiently adapt SAM to the dataset labeled with affordance maps instead of its original output masks, we introduce this module. Similar to BLIP-2 [26] and DETR [8], we introduce a set of learnable queries termed affordance queries. These queries, denoted as $\mathbf{Q}_{a}\in{N_{s}\times D_{m}}$, first repeat over batch dimension and conduct cross-attention with the text features that interact with the image in the multimodal encoder. we expect these queries to extract information concerning the affordance function from these semantic features. What’s more, since different layers of SAM’s features often exhibit different levels of granularity and knowledge [17], and affordance may correspond to multiple parts of an object. We consider a diverse set of granularities can be beneficial. Therefore, the second cross-attention operation is performed between the processed affordance queries and fused visual features of SAM’s multiple layers formulated as:
@@ 3.2 Affordance-Adaption Module
3.3 Coarse-to-Fine Training Recipe
In order to fully leverage SAM’s generalization capacity for affordance grounding, we elaborately crafted a coarse-to-fine training recipe shown in Figure 2(b), which consists of three stages. The overall training data is displayed in Table 1. In all stages, we use a template of “<affordance function> <object_name>”, for instance, “wear hat”, as the text prompt for AffordanceSAM. The detailed dataset setup are provided in Appendix 2.
Training stage1. In most training datasets for multimodal models and text-prompted segmentation models, such as LAION-400M [44], RefCLEF [16], and RefCOCO [50], there are very few objects with affordance properties and very little textual content that contains verbs. However, it is crucial for an affordance grounding model to familiarize itself with such objects and understand affordance verb information for locating potential “action-possibilities” regions. Based on the above observations, we decided to let EVF-SAM first meet this requirement from a coarse perspective without any change in output form. Therefore, we select three datasets, PADv2 [51], Handal [13], and RGB-D Part Affordance [33], which satisfy both objects contained with affordance functions and annotations of masks. We then combine them according to the categories of objects and affordance functions. We only use EVF-SAM baseline without the affordance-adaption module and fine-tune multimodal encoder and prompt encoder in this stage. The training loss function is given by
@@ 3.3 Coarse-to-Fine Training Recipe
where the overall loss function $\mathcal{L}$ is the weighted sum of DICE loss $\mathcal{L}_{dice}$ [45] and BCE loss $\mathcal{L}_{bce}$ [43], determined by $\lambda_{dice}$ and $\lambda_{bce}$.
@@
Cross-View-AG [29] CVPR’22 1.787 0.285 0.829 2.092 0.209 0.138
LOCATE [24] CVPR’23 1.405 0.372 1.157 1.829 0.282 0.276
AffordanceLLM [39] CVPR’24 1.463 0.377 1.070 1.661 0.361 0.947
OOAL [25] CVPR’24 1.070 0.461 1.503 1.302 0.410 1.119
AffordanceSAMSAM1 this work 1.083 0.543 1.800 1.128 0.514 1.761
AffordanceSAMSAM2 this work 1.141 0.530 1.631 1.214 0.496 1.536
Table 2: Quantitative results on the Easy split and Hard Split of AGD20K. The best and second-best results are highlighted in bold and underlined. SAM1 and SAM2 denote different versions of SAM backbones (SAM-ViT-H [20] and SAM2-hiera-L [41]) for initialization.
Algorithm 1 Affordance maps post-process algorithm in stage2.
Input: Affordance map $M_{al}$ generated by LOCATE and threshold $\gamma$.
1:  $\gamma_{1}=\gamma\cdot max(M_{al})$
2:  $\gamma_{2}=0.4\cdot\gamma_{1}$
3:  $\gamma_{3}=0.2\cdot\gamma_{2}$
4:  $M_{ap}=where(M_{al}\geq\gamma_{1},M_{al},\frac{{M_{al}}^{2}}{\gamma_{1}})$
5:  $M_{ap}=where(M_{ap}\geq\gamma_{2},M_{ap},\frac{{M_{ap}}^{2}}{\gamma_{2}})$
6:  $M_{ap}=where(M_{ap}\geq\gamma_{3},M_{ap},0)$
Output: Processed affordance map $M_{ap}$.
Training stage2. As mentioned in Section 1 and Table 1, images with dense pixel labeling affordance maps are scarce and it can be costly for us to annotate. In practice, we find that training only on the hundreds of labeled images could not fully convert SAM’s output mask into an affordance map. Hence, we utilize the state-of-the-art weakly supervised model on AGD20K, LOCATE [24], to label the vast majority of the remaining unlabeled images in AGD20K. However, The raw output affordance heatmaps of LOCATE are usually suboptimal, which exist a large area of low thermal regions unrelated to the target. Hence, we design a post-processing program showed in Algorithm 1 to mitigate such issue. These processed pseudo labels help bridge the gap between binary segmentation masks from stage1 and the detailed affordance heatmaps needed for the final output, serving as an intermediate step in our coarse-to-fine training pipeline. During training, we add affordance-adaption module and keep only SAM image encoder frozen. we follow AffordanceLLM [39] to use binary focal loss [27] ($\mathcal{L}_{focal}$) and set the weight of positive examples to be 0.9 and that of negative ones to be 0.1, as there are often more negatives than positives in an affordance map.
Training stage3. In this stage, we use the high-quality human-labeled samples in AGD20K to supervise the model. We aim at regulating the model to generate the most precise and fine-grained affordance maps. We keep the loss function and trainable modules unchanged compared to stage2.
4 Experiments
4.1 Experimental Setup
Implementation details. In stage1, we initialize our AffordanceSAM with EVF-SAM [52] and train for 13 epochs with base learning as 0.00002. In stage2, we use the weights obtained from the first stage and randomly initialize affordance-adaption module, we train the model with the same number of epochs and base learning as stage1. In stage3, we further train our model for 26 epochs with larger base learning as 0.00004. In all training stages, we employ an AdamW optimizer [19] with a cosine learning rate scheduler, and the total batchsize per iteration is 32 using 4 NVIDIA A100 (80GB) GPUs.
Evaluation. We follow AffordanceLLM [39] to have two splits (easy split and hard split) of AGD20K to test our model. The easy split is the original unseen split of AGD20K, which has a lot of similarities between the objects in the train and test sets. The hard split ensures that there is no overlap between the object categories in the train and test set (both in our stage 2 and stage 3). Meanwhile, we follow the metrics provided in AGD20K to evaluate our model, which is Kullback-Leibler Divergence (KLD) [7], Similarity (SIM) [46] and Normalized Scanpath Saliency (NSS) [37]. For KLD, the lower the better, and for SIM and NSS, the higher, the better.
Methods for comparison. As we indicate in Section 1, affordance grounding methods can be mainly summarized into two main categories: weakly supervised and fully supervised methods. We compare our approach against representative weakly supervised baselines (Cross-View-AG [29] and LOCATE [24]), and fully supervised baselines (AffordanceLLM [39] and OOAL [25]). we do not conduct the qualitative comparison with AffordanceSAM because it is not open-source and lacks sufficient implementation details to be reproduced.
We also provide more details about our implementation including training hyperparameters and model configuration, details of each metric, and detailed descriptions of each baseline method for comparison in Appendix  3$\sim$5.
4.2 Main Results
@@
Figure 3: Qualitative comparison with other methods on AGD20K dataset. Cross-View-AG and LOCATE often make affordance predictions with a large area of low thermal regions unrelated to the target. On the contrary, OOAL focuses on a small area but sometimes it’s completely wrong. Our AffordanceSAM can precisely recognize the part of the object that contains the affordance function.
Quantitative Results on AGD20K. Table 2 presents the quantitative results of our proposed AffordanceSAM alongside other baselines on both the easy split and hard split of the AGD20K. Our AffordanceSAM , particularly the SAM1 variant, demonstrates superior performance across multiple metrics. On the easy split, it achieves the highest scores in SIM (0.543) and NSS (1.800). Most importantly, the advantages of our approach become more pronounced on the challenging hard split, where the objects in the test set bear little to none resemblance to the ones in the training set. To be specific, both two versions of AffordanceSAM achieve superior performance, and our best version reduces the KLD by 17.4%, while simultaneously increasing the SIM and NSS scores by 10.4% and 64.2% respectively when compared to the previous best performing method OOAL [25]. This delightful result suggests that our method successfully inherits SAM’s strong generalization ability and generalizes well to more challenging scenarios.
Qualitative Results on AGD20K. As shown in Figure 3, we present four examples from the validation set of AGD20K to see how AffordanceSAM outperforms other methods. For instance, when handling the object “cup”, AffordanceSAM demonstrates a superior understanding of affordance functions compared to other methods and generates more accurate and precise affordance maps according to different actions “hold” and “drink”. Furthermore, benefiting from SAM’s ability to accurately segment objects at the pixel level, our AffordanceSAM can sometimes give even more precise affordance maps than the ground-truth (see the middle two lines of Figure 3).

### v2

@@ Abstract
Building a generalized affordance grounding model to identify actionable regions on objects is vital for real-world applications. Existing methods to train the model can be divided into weakly and fully supervised ways. However, the former method requires a complex training framework design and can not infer new actions without an auxiliary prior. While the latter often struggle with limited annotated data and components trained from scratch despite being simpler. This study focuses on fully supervised affordance grounding and overcomes its limitations by proposing AffordanceSAM, which extends SAM’s generalization capacity in segmentation to affordance grounding. Specifically, we design an affordance-adaption module and curate a coarse-to-fine annotated dataset called C2F-Aff to thoroughly transfer SAM’s robust performance to affordance in a three-stage training manner. Experimental results confirm that AffordanceSAM achieves state- of-the-art (SOTA) performance on the AGD20K benchmark and exhibits strong generalized capacity.
Introduction
@@
Affordance grounding (Gibson 2014) refers to finding potential “action possibilities” regions of an object, which plays a key role in bridging the gap between visual perception and robotic action. Recently, attempts made to endow models to have such grounding abilities can be broadly divided into two ways. The first type of methods is weakly supervised affordance grounding (Li et al. 2023a; Xu and Mu 2025; Wang et al. 2025), which aims to identify affordance regions on objects using human-object interaction images and egocentric object images without dense labels. The affordance maps are obtained via class activation mapping (CAM) (Zhou et al. 2016) or an auxiliary prior model (Kirillov et al. 2023). However, these methods require a complex training framework design since there are two branches (exocentric and egocentric) that need to be balanced and optimized during training. Moreover, the generalization performance of these models is suboptimal, as their supported classes are fixed (e.g, LOCATE (Li et al. 2023a)) or very few parameters are optimized on the affordance task (e.g., PLSP (Xu and Mu 2025)). To tackle this problem, the second type of methods uses affordance maps to directly supervise the models (Qian et al. 2024; Li et al. 2024). However, some important affordance components trained from scratch (e.g., decoder) on only hundreds of pieces of manually labeled training data make it insufficient to obtain a highly generalized model. Thus, it is important for the fully supervised affordance grounding family to find a strong foundation model that is naturally suitable for the affordance grounding task and scale up the supervised data to obtain the ideal generalized model.
Figure 1: Performance Comparison: The circle area indicates the number of training data, with better-performing models positioned toward the upper right. Our AffordanceSAM and C2F-Aff data can respectively serve as an excellent base model and training data. Integrating the two achieves a performance far ahead of other candidates.
In this work, we try to decompose and address this problem in a step-by-step manner:
@@
(i) Incorporating a suitable and generalized vision foundation model. With the recent advancements in large-scale pre-trained foundation models, many works have attempted to leverage the prior knowledge for downstream task transfer (Wang et al. 2023a; Han and Lim 2024). We believe this paradigm is also effective for the affordance grounding task. Given that affordance grounding inherently requires pixel-level localization, we explore the possibility of incorporating SAM (Segment Anything Model) (Kirillov et al. 2023; Ravi et al. 2024) since SAM not only demonstrates remarkable generalization capabilities but also excels in precise object localization after pre-trained on large-scale masked-labeled data. Moreover, the model structure and decoder output format of segmentation and affordance grounding are highly similar. Considering these similarities between segmentation and affordance grounding, we hypothesize that incorporating SAM could be well-suited for our problem.
Figure 2: Differences of two task, where segmentation focuses on separating objects according to the prompt, but affordance emphasizes grounding the possible effective part of objects based on the affordance query.
@@
Although there are some similarities between segmentation and affordance grounding, task differences still exist. As shown in Figure 2, firstly, affordance grounding requires the recognition of functional parts of objects based on verb queries, while segmentation requires the separation of objects; the ground-truth of segmentation and affordance also differs slightly: the former uses a binary mask, while the latter uses a heatmap that encodes functional possibilities.
@@
(ii) Effectively transferring the capabilities of SAM through data and model structure aspects. As mentioned above, there exist some differences of these two tasks, to fully unlock SAM’s capacity for affordance grounding, we first design an affordance-adaption module that utilizes learnable queries to interact with text and image features. These queries are then sent into the decoder to refine the original mask produced by SAM. We hope that this prompt learning-like module can effectively search for affordance-relevant parts in the comprehensive knowledge of SAM. We then scale up the fully supervised data in a coarse-to-fine annotated manner, which we named C2F-Aff. The dataset consists of three parts with each corresponding to a specific label format, and thus the training process is divided into three stages. In the first data part and training stage, we curate a mask-labeled dataset by merging and reclassifying existing datasets, focusing on object-affordance pairings. This enables the model to learn basic object-verb relationships. After that, we use a weakly supervised model to generate pseudo-labels for unannotated images in AGD20K (Luo et al. 2022) and post-process them to obtain annotations with higher quality to finetune the model. Finally, we use the high-quality, human-labeled data in AGD20K for further fine-tuning the model to boost the performance.
In our comprehensive experiments, we show that AffordanceSAM and C2F-Aff data can severally serve as an excellent base model and training data, which is evidenced by the leading performance of AffordanceSAM against other methods using the same amount of data and the performance enhancement of a wide range of fully supervised methods once trained on our C2F-Aff data.
@@
  • Showing that SAM is natural fit for affordance grounding with analysis and proposed AffordanceSAM.
  • Introducing C2F-Aff dataset and a three-stage training scheme that boosts the performance of both AffordanceSAM and prior fully-supervised methods.
  • Achieving SOTA performance on the AGD20K benchmark, as well as showing the evidence to generalize to novel actions and objects.
Related Work
We discuss the most relevant studies here and provide a more discussion in Appendix 1.
Affordance Grounding. Unlike detection (Zhao et al. 2019) and segmentation (Minaee et al. 2021) that yield bounding boxes or binary masks of the desired objects, affordance grounding (Gibson 2014) needs a model to output the functional possibilities in a object. Early works (Nagarajan, Feichtenhofer, and Grauman 2019; Mai, Yang, and Luo 2020; Pan et al. 2021) have tried to equip the model with such capacity. But because of the data restriction at that time, these models can only recognize a few types of objects and actions. Next, Luo et al. collect the first large-scale affordance dataset called AGD20K (Luo et al. 2022) and annotated a few of the data with affordance maps (1675/23816). After that, many works engined by AGD20K have emerged. For example, LOCATE (Li et al. 2023a) localizes and extracts affordance-specific information in exocentric and transfers this knowledge to egocentric in a weakly supervised manner; AffordanceLLM (Qian et al. 2024) builds a vision-language learning framework based on LLaVA (Liu et al. 2023) to conduct the open affordance learning. PLSP (Xu and Mu 2025) uses an additional semantic prior (e.g., SAM (Kirillov et al. 2023)) [^1] While both PLSP and our work use SAM, PLSP only uses SAM as a prior and refiner to guide and assist the main model branch. By contrast, our method directly tunes SAM to be an affordance grounding model in a fully supervised manner. to guide the traditional two-branch weakly supervised training process. However, none of the above works show satisfactory generalization ability to recognize regions for unseen objects and affordance actions (reasons are explained in the Introduction, and evidence is shown in our experiment results). In this study, we aim to find a foundation model that is naturally suitable for the task and scale up the annotated data to obtain such an ideal model.
@@
Figure 3: (a). Architecture: AffordanceSAM is built upon EVF-SAM  (Zhang et al. 2024) with an additional affordance-adaption module. (b). Training: Trained on purposed C2F-Aff dataset, AffordanceSAM adopts a coarse-to-fine training recipe.
Foundation Model for Downstream Tasks. Vision foundation models (e.g., SAM (Kirillov et al. 2023), CLIP (Radford et al. 2021), DINOv2 (Oquab et al. 2023)) are large-sized models trained over vast amounts of data . After training, they are suitable as a starting point for a variety of downstream tasks under limited data conditions. Recently, many works have used different foundation models for different downstream tasks (Zhao et al. 2023a; Wang et al. 2023a; Han and Lim 2024; Khattak et al. 2023). For example, ActionCLIP (Wang et al. 2023a) proposes a “pre-train, prompt and fine-tune” strategy to transfer CLIP to action recognition; FM-FSOD (Han and Lim 2024) uses DINOv2 to extract fine-grained features for few-shot object detection. Our work also shares similarities, while we focus on adapting SAM to the affordance grounding task with proposed affordance-adaption module and curated data.
Approach
The overall model architecture and training procedure of AffordanceSAM are shown in Figure 3(a) and Figure 3(b). As our model consists of two primary components: the EVF-SAM (Zhang et al. 2024) [^2] EVF-SAM can be seen as an extension version of SAM that support text prompts. In this paper, we use term ‘EVF-SAM’ to refer the specific model part and ‘SAM’ for our general argument. baseline and the affordance-adaption module, we begin with describing each, and finaly, we provide a comprehensive elucidation of our coarse-to-fine affordance (C2F-Aff) dataset and training recipe.
Preliminary
Part Data Source Label Type Num.
@@ 3 Labeled Part of AGD20K Human Annotated Affordance Map 1135 (Easy Split)
Table 1: Overview of the training dataset C2F-Aff. This table summarizes the data source, format of groundtruth, the number of samples for training in each stage. More details can be found in Appendix 2.
Baseline EVF-SAM. As we need to provide text prompts that contain affordance queries to the model but the original SAM lacks language understanding abilities. We choose to start from EVF-SAM (Zhang et al. 2024) which is the SOTA method among attempts (Zhang et al. 2024; Lai et al. 2024; Ren et al. 2024; Zhao et al. 2023b) that empowered SAM to follow text instructions. We seek to leverage its robust segmentation capabilities for affordance grounding.
The EVF-SAM original framework incorporates four key components: a multimodal encoder (BEIT-3 (Wang et al. 2023b)) denoted as $\mathcal{E}_{M}$, a prompt encoder $\mathcal{E}_{P}$, a SAM image encoder $\mathcal{E}_{I}$, and a SAM mask decoder $\mathcal{D}$. The input image is firstly processed through two parallel paths: (i): conversion to SAM image tokens $\mathbf{I}_{s}\in\mathbb{R}^{B\times N_{s}\times D_{s}}$, and (ii): transformation to multimodal image tokens $\mathbf{I}_{m}\in\mathbb{R}^{B\times N_{m}\times D_{m}}$, where $B$ represents batch size, $N$ sequence length, and $D$ feature dimension. Meanwhile, the text input is first tokenized to produce text tokens $\mathbf{T}\in\mathbb{R}^{B\times N_{t}\times D_{m}}$. These text and multimodal image tokens are then combined with a learnable [CLS] token $\in\mathbb{R}^{B\times 1\times D_{m}}$ and processed by the multimodal encoder:
@@
Affordance-Adaption Module
As shown in Figure 2, there exists a substantial disparity between the output of the segmentation task and the affordance grounding task. To efficiently adapt SAM to the dataset labeled with affordance maps instead of its original output masks, we introduce this module. Similar to BLIP-2 (Li et al. 2023b) and DETR (Carion et al. 2020), we introduce a set of learnable queries termed affordance queries. These queries, denoted as $\mathbf{Q}_{a}\in{N_{s}\times D_{m}}$, first repeat over batch dimension and conduct cross-attention with the text features that interact with the image in the multimodal encoder. we expect these queries to extract information concerning the affordance action from these semantic features. What’s more, since different layers of SAM’s features often exhibit different levels of granularity and knowledge (Ke et al. 2024), and affordance may correspond to multiple parts of an object. We consider a diverse set of granularities can be beneficial. Therefore, the second cross-attention operation is performed between the processed affordance queries and fused visual features of SAM’s multiple layers formulated as:
@@
Coarse-to-Fine Affordance Dataset and Training
To fully leverage SAM’s generalization capacity for affordance grounding, we craft a coarse-to-fine training dataset (C2F-Aff) and training recipe shown in Figure 3(b). The brief summary of C2F-Aff is displayed in Table 1, and more information on our data collection, division, organization and cleaning can be found in Appendix 2. We divide the data and training into three parts based on the form of the labels, while we always use a template of “<affordance action> <object_name>”, for instance, “wear hat”, as the text prompt for AffordanceSAM.
Dataset Part and Training Stage 1. In most training datasets for multimodal models and text-prompted segmentation models, such as LAION-400M (Schuhmann et al. 2021), RefCLEF (Kazemzadeh et al. 2014), and RefCOCO (Yu et al. 2016), there are very few objects with affordance properties and very little textual content that contains verbs. However, when transferring a foundation model to affordance grounding task in a fully supervised manner (e.g, DINOv2 in OOAL, LLaVA in AffordanceLLM, and SAM in our method), it is crucial for them to familiarize themselves with such objects and understand the affordance verb information for locating potential “action-possibility” regions. Thus, In the first part of our C2F-Aff dataset, we select three datasets, PADv2 (Zhai et al. 2022), Handal (Guo et al. 2023), and RGB-D Part Affordance (Myers et al. 2015), which satisfy both objects contained with affordance actions and annotations of masks. We then combine them according to the categories of objects and affordance actions (details are in Appendix 2). And for We only use EVF-SAM baseline without the affordance-adaption module and fine-tune multimodal encoder and prompt encoder in this stage. The training loss function is given by
@@
where the overall loss $\mathcal{L}$ is the weighted sum of DICE loss $\mathcal{L}_{dice}$ (Sudre et al. 2017) and BCE loss $\mathcal{L}_{bce}$ (Ruby and Yendapalli 2020), determined by $\lambda_{dice}$ and $\lambda_{bce}$.
Dataset Part and Training Stage 2. As mentioned in the Introduction, images with dense pixel labeling affordance maps are scarce and it can be costly for us to annotate. In practice, we find that training only on the hundreds of labeled images could not fully unlock SAM’s capacity. Hence, we utilize a weakly supervised model on AGD20K, LOCATE (Li et al. 2023a), to label the vast majority of the remaining unlabeled images in AGD20K. However, The raw output affordance heatmaps of LOCATE are usually suboptimal, which exist a large area of low thermal regions unrelated to the target. Hence, we design a post-processing program showed in Algorithm 1 to mitigate such issue. These processed pseudo labels help bridge the gap between binary segmentation masks from part 1 and the detailed affordance heatmaps needed for the final output, serving as an intermediate step in our C2F-Aff dataset. During training, we add affordance-adaption module and keep only SAM image encoder frozen. we follow AffordanceLLM (Qian et al. 2024) to use binary focal loss (Lin 2017) ($\mathcal{L}_{focal}$) and set the weight of positive examples to be 0.9 and that of negative ones to be 0.1, as there are often more negatives than positives in an affordance map.
Algorithm 1 Affordance maps post-process algorithm.
Input: Affordance map $M_{al}$ generated by LOCATE and threshold $\gamma$.
[…]
Output: Processed affordance map $M_{ap}$.
Dataset Part and Training Stage 3. To achieve the final high-quality supervised fine-tuning for boosting performance, we use the high-quality human-labeled samples in AGD20K as the third part of our C2F-Aff dataset and use them to supervise the model. We aim at regulating the model to generate the most precise and fine-grained affordance maps. We keep the loss function and trainable modules unchanged compared to stage 2.
Table 2: Quantitative results on AGD20K benchmark. The best and second-best results are marked in bold and underlined.
@@
Weakly supervised methods trained on AGD20K-Weakly ($\sim$12K)
Cross-View-AG (Luo et al. 2022) CVPR’22 1.787 0.285 0.829 2.092 0.209 0.138
LOCATE (Li et al. 2023a) CVPR’23 1.405 0.372 1.157 1.829 0.282 0.276
R-Mamba (Wang et al. 2025) CVPR’25 1.310 0.397 1.279 - - -
PLSP (Xu and Mu 2025) ICLR’25 1.153 0.437 1.418 1.401 0.395 1.109
Fully supervised methods trained on AGD20K-Fully ($\sim$1K)
AffordanceLLM (Qian et al. 2024) CVPR’24 1.463 0.377 1.070 1.661 0.361 0.947
OOAL (Li et al. 2024) CVPR’24 1.070 0.461 1.503 1.302 0.410 1.119
AffordanceSAM this work 1.271 0.486 1.597 1.327 0.423 1.502
Fully supervised methods trained on C2F-Aff ($\sim$40K + $\sim$12K + $\sim$1K)
AffordanceLLM (Qian et al. 2024) CVPR’24 1.170 0.482 1.425 1.312 0.405 1.293
OOAL (Li et al. 2024) CVPR’24 0.974 0.504 1.650 1.119 0.442 1.364
AffordanceSAM this work 1.083 0.543 1.800 1.128 0.514 1.761
@@
Figure 4: Qualitative comparison on AGD20K dataset. Cross-View-AG and LOCATE often make affordance predictions with a large area of low thermal regions unrelated to the target. On the contrary, OOAL focuses on a small area but sometimes it’s completely wrong. Our AffordanceSAM can precisely recognize the part of the object that contains the affordance action.